if P(notA)=0.5, P(notB)= 0.4, P(A and B)=0.3, find P(A or B).
Answer by robertb(5830)

P(A') = 0.5 ===> P(A) = 0.5
P(B') = 0.4 ===> P(B) = 0.6
===> P(A∪B) = p(A) + p(B) - P(A∩B) = 0.5 + 0.6 - 0.3 = 0.8.